# Reborn: Reinforcement-Learned Boundary Segmentation with Iterative Training for Unsupervised ASR

### v2

@@ 1 Introduction
Automatic speech recognition (ASR) systems convert speech signals into their transcription texts. Most state-of-the-art (SoTA) ASR systems are supervisedly trained with a large amount of labeled speech-text paired data [9, 49]. However, more than 80% of languages in the world have limited access to speech-text paired data [48], and collecting such paired data requires intensive labor and costs, limiting the applicability of supervised ASR systems to these languages. This leads to significant efforts devoted to developing unsupervised ASR (UASR) systems, which aim to learn the mapping between speech signals and textual transcriptions (words or phonemes) without any speech-text pairs [31, 10, 4, 28, 17].
UASR models learn to align the distribution of input speech signal and output text without paired data. Learning to match two distributions of sequences unsupervisedly has been extensively studied in unsupervised neural machine translation (NMT) [1], where the aim is to learn a neural network that can translate text in a source language to text in a target language without paired data [24, 25]. Prior works show that adversarial training [19] can be used to learn such mappings [24, 25], which employs a generator learning to translate and a discriminator learning to distinguish generated text from the real text in the target language.
@@
This paper focuses on learning better segmental boundaries to improve the mapping between segmented speech and textual transcription. We propose Reborn (Reinforcement-Learned Boundary Segmentation with Iterative Training), an unsupervised ASR framework with a segmentation model and a phoneme prediction model. The segmentation model determines segmental structure boundaries in speech signals, while the phoneme prediction model assigns a phoneme to each segmental structure. After properly initializing the phoneme prediction model, we use an iterative algorithm that alternates between two stages to train the segmentation model and refine the phoneme prediction model: The first stage trains the segmentation model. Since we do not have the ground truth segmental boundaries, we use reinforcement learning to train the segmentation model to favor segmentations that yield phoneme sequence predictions with a lower perplexity. We experiment with various learning objectives and implement them as reward functions for learning the segmentation model. In the second stage, based on the boundaries predicted by the segmentation model learned in the previous stage, we use an adversarial loss similar to generative adversarial networks (GANs) [19] to train the phoneme prediction model.
@@
We conduct extensive experiments on LibriSpeech [42], TIMIT [18], and Multilingual LibriSpeech (MLS) [47] to compare the phoneme/phone error rate (PER) and word error rate (WER) with prior works. We show that our method outperforms all prior UASR methods on LibriSpeech, TIMIT, and five languages in MLS when using the same amount of training data. We perform thorough ablation studies to show that iterative training and the rewards we design are critical to the performance of Reborn. By analyzing the segmental structure obtained by our segmentation model, we find that the segmental structures are acoustic units smaller than phonemes, which helps the phoneme prediction model predict more accurate transcriptions.
2 Related Work: Unsupervised ASR
@@
An ASR model takes speech signals as the input and predicts the textual transcriptions. Unsupervised ASR (UASR) aims to train an ASR model without access to paired speech-text data. Instead, the only available training data is unlabeled speech data and unlabeled text data, while the correspondence between speech and text is not available. In this paper, we follow prior UASR works [31, 4] to predict phoneme transcriptions from speech signals. Learning the mapping between speech signals and phoneme sequences can be formulated as a distribution matching problem, where we want to learn an ASR model whose output phoneme sequences match the distribution of real phoneme sequences.
@@
Yeh et al. [60] solve this distribution matching problem by an unsupervised loss function based on the empirical output distribution matching (ODM) [32], which guides the ASR model to output phoneme sequences that have statistic distributions close to real phoneme sequences.
@@
Liu et al. [31] propose to use the generative adversarial network (GAN) [19] for UASR. GAN solves the distribution matching problem by training a generator whose output distribution resembles a target distribution, and a discriminator is trained to distinguish the outputs of the generator and the samples from the target distribution. When using GAN to train unsupervised ASR, the generator is the phoneme prediction model. The phoneme prediction model takes in speech features of the segmental structures in the speech signal, and outputs a phoneme transcription. Prior works rely on hand-crafted or separately trained unsupervised phoneme segmentation models to find the boundary of the segmental structures [57, 31, 10].
@@
Figure 1: (a) and (b): Reborn iterates between using RL to train the segmentation model and using adversarial training to train the phoneme prediction model. (c): An illustration of the segmentation/boundary merging. Given the original segmentation and the predicted phoneme sequence, we merge the segments that result in the same phoneme prediction into the same segment, yielding the merged boundary.
@@
In these prior methods, the segmentation is either non-trainable or learned separately from the phoneme prediction model. Reborn differs from prior works by introducing a trainable segmentation model tailored for the phoneme prediction model, and the segmentation model and phoneme prediction model can be iteratively polished. wav2vec-U 2.0 [28] simplifies the feature preprocessing step in wav2vec-U and does not explicitly consider the segmentation structure in their model. Reborn has performance better than wav2vec-U 2.0 on almost all datasets, showing that learning the segmentation boundary for feature preprocessing is important to the performance of UASR.
Although originated to be a cross-modality distribution matching problem, recent UASR is also related to representation learning. For example, the speech features might be extracted from self-supervised speech foundation models; and the segmentation problem is also commonly investigated in tasks like acoustic unit discovery. We leave the discussions about the two related topics in Appendix F.
@@ 3 Method: Reborn
The difficulty of mapping speech signals to their corresponding phoneme transcriptions lies in the segmental structures in speech whose boundaries are unknown. To tackle this challenge, we propose Reborn, which trains a UASR system with unpaired speech and text data. Reborn contains a segmentation model and a phoneme prediction model. The segmentation model takes the speech feature as input and determines the boundaries of the segmental structures in the speech signal, and the phoneme prediction model predicts a phoneme for each segmental structure. We will use the terms segment and segmental structure interchangeably. In our paper, we do not use the term segment in the exact sense as in linguistics, where segment refers to discrete units that can be identified in the stream of speech [14]. We use segment to broadly refer to a span in the speech, which may be a meaningful unit (e.g., a phone or word) or a span in the speech identified by the segmentation model (Section 3.1).
The overall training process of Reborn is outlined in Figure 1. First, we initialize the phoneme prediction model using wav2vec-U (Section 3.3). Next, the training of Reborn iterates through two stages: Stage 1 (Figure 1(a), Section 3.1): Training the segmentation model to learn a better segmentation until the segmentation model converges while fixing the phoneme prediction model. Stage 2 (Figure 1(b), Section 3.2): Training the phoneme predictor based on the segment boundaries predicted by the segmentation model until the phoneme prediction model converges. The iterative process ceases when the UASR performance does not improve over the previous iteration. In UASR, we cannot use PER as a validation metric to select the best model or determine when to stop training. We use an unsupervised evaluation metric to select the best model, detailed in Appendix E.4.
@@ 3.1.1 Segmentation Model
Given an input speech signal, we use a self-supervised feature extractor model to extract the speech features $X=[x_{1},\cdots,x_{T}]$ from the waveform. The segmentation model takes the speech features and predicts the boundary of the segmental structure in the speech features. For each feature $x_{i}$, the segmentation model assigns a binary value $\hat{b}_{i}$, indicating whether $x_{i}$ is the first frame of a segment. We split $X$ into segments based on $\hat{b}=\hat{b}_{1:T}$ and mean-pool the features within the same segment. The mean-pooled features $[s_{1},\cdots,s_{T^{\prime}}]$ are forwarded through the phoneme prediction model to obtain the phoneme prediction $[y_{1},\cdots,y_{T^{\prime}}]$. The resulting phoneme prediction will be de-duplicated into $[y^{\prime}_{1},\cdots,y^{\prime}_{M}]$ by removing the same consecutive phoneme.
@@ 3.1.2 Training the Segmentation Model with RL
To help the prediction of the phoneme prediction model, we want to train the segmentation model to capture the segmental structure in speech. However, the optimal segmentation boundary is not available for training. Recognizing that the segmentation quality directly affects the phoneme prediction of the phoneme prediction model, we estimate the quality of the phoneme prediction to guide the segmentation model training. The phoneme prediction model is from the previous Reborn iteration or initialized from wav2vec-U in the first Reborn iteration, and it is fixed when training the segmentation model. Since the effect of feature segmentation on phoneme prediction is non-differentiable, we employ RL to train the segmentation model. The estimated quality of the phoneme sequence serves as the RL reward.
We train the segmentation model using the policy gradient method [54] based on REINFORCE algorithm [59]. For each utterance, we calculate an utterance-wise reward $R$ (defined in Eq. 4 in the next subsection) from the de-duplicated phoneme prediction to train the segmentation model. Based on the policy gradient method, the segmentation model $\pi_{\theta}$ is optimized using the following gradient: $\mathbb{E}_{\hat{b}\sim\pi_{\theta}}\left[\nabla_{\theta}\log\pi_{\theta}(\hat{b}|X)R\right]$, where $\mathbb{E}$ is taken over $\hat{b}$ sampled from $\pi_{\theta}$ and approximated with the mean of a batch of training sequences.
@@ 3.1.3 Reward Designs
We introduce some notations: Given an utterance, we use the currently trained segmentation model $\pi_{\theta}$ for segmentation and the phoneme prediction model trained in the last Reborn iteration to obtain a deduplicated phoneme prediction sequence, denoted by $Y^{\prime}_{\theta}$. For the same utterance, we can use the segmentation model trained in the previous Reborn iteration $\pi_{\theta-1}$ for features segmentation and the same phoneme prediction model to obtain another de-duplicated phoneme prediction sequence, denoted as $Y^{\prime}_{\theta-1}$. In the first Reborn iteration, $Y^{\prime}_{\theta-1}$ is the de-duplicated phoneme prediction from wav2vec-U.
Perplexity Difference Reward
The perplexity difference reward is designed to favor phoneme segmentation better than the segmentation learned in the previous iteration. Intuitively, a better phoneme segmentation prediction $\hat{b}_{1:T}$ should yield a more reasonable phoneme prediction $y^{\prime}_{1:M}$. We use perplexity (PPL), the negative likelihood of a phoneme sequence scored by a phoneme language model (LM), to evaluate how reasonable a phoneme sequence is. Perplexity measures how likely a phoneme sequence can be observed in the real world; a phoneme sequence with a lower perplexity means it is more likely to be observed in real-world phoneme datasets. We use a 4-gram phoneme LM trained on the phonemicized unlabeled text corpora. To guide the segmentation model to generate a segmentation with a better phoneme prediction than the phoneme predictions obtained in the previous iteration, we define perplexity difference reward as follows:
@@
Edit Distance Reward
We use Levenshtein distance $d_{\text{Lev}}$ as the edit distance to calculate the difference between $Y^{\prime}_{\theta-1}$ and $Y^{\prime}_{\theta}$. We take the normalized negative edit distance as the reward, which is defined in Eq. 2
Length Difference Reward
We use the length difference reward $R_{\text{len}}$ to guide the segmentation model to predict segmentation such that the length of $Y^{\prime}_{\theta}$ does not differ significantly from the length of $Y^{\prime}_{\theta-1}$. The length difference reward for an utterance is defined in Eq. 3, where $|Y^{\prime}|$ is the length of $Y^{\prime}$.
@@
where $c_{\text{ppl}}$, $c_{\text{edit}}$ and $c_{\text{len}}$ are the weighting coefficients. During training, $R_{\text{ppl}}$, $R_{\text{edit}}$ and $R_{\text{len}}$ are normalized within each training batch. Appendix E.4 discusses how the coefficients are determined without speech-text paired data.
@@ 3.1.4 Initializing $\pi_{\theta}$ with Behavior Cloning
Before training the segmentation model $\pi_{\theta}$ with RL, we initialize it with behavior cloning (BC) [44, 45]. BC uses the supervised objective to train the segmentation model to predict the boundaries in speech features. Given speech features $x_{1:T}$, the segmentation model is trained to predict the 0/1 boundary $\hat{b}_{i}$ by using some boundaries $b_{i}$ as the ground truth. The ground truth boundary is the merged boundary (Section 3.1.5) obtained using $\pi_{\theta-1}$, the segmentation model learned in the previous iteration. The BC process is a binary classification task that can be trained with cross-entropy loss. In the first Reborn iteration, we use the $k$-means-based segmentation boundary from wav2vec-U as the BC target.
@@ 3.2 Stage 2: Training the Phoneme Prediction Model
After obtaining the merged phoneme boundaries, we train the phoneme predictor model using GAN training. The phoneme prediction model is the generator in GAN training. The generator aims to output phoneme predictions that look like real phoneme sequences to fool the discriminator. The discriminator takes in a phoneme sequence, which can be the output of the generator or a phoneme sequence from the unpaired text corpora, and the goal of the discriminator is to distinguish whether the input phoneme sequences are outputs of the generator. The generator and discriminator are updated using the loss in GAN training. In this stage, the parameters of the segmentation model are not updated, and the generator is initialized from the previous iteration. We discuss the effect of applying boundary merging and parameter initialization in Appendix C.3

### v3

@@ 1 Introduction
Automatic speech recognition (ASR) systems convert speech signals into their transcription texts. Most state-of-the-art ASR systems are trained with dense supervision using a large amount of labeled speech-text paired data [10, 52]. However, more than 80% of languages in the world have limited access to speech-text paired data [51], and collecting such paired data requires intensive labor and costs, fundamentally limiting the applicability of supervised ASR systems to these languages. This leads to significant efforts devoted to developing unsupervised ASR (UASR) systems, which aim to learn the mapping between speech signals and textual transcriptions (words or phonemes) without any speech-text pairs [33, 11, 4, 30, 18].
UASR models learn to align the distribution of input speech signal and output text without paired data. Learning to match two distributions of sequences unsupervisedly has been extensively studied in unsupervised neural machine translation (NMT) [1], where the aim is to learn a neural network that can translate text in a source language to text in a target language without paired data [26, 27]. Prior works show that adversarial training [20] can be used to learn such mappings [26, 27], which employs a generator learning to translate, and a discriminator learning to distinguish generated text from the real text in the target language.
@@
This paper focuses on learning better segmental boundaries to improve the mapping between segmented speech and textual transcription. We propose Reborn (Reinforcement-Learned Boundary Segmentation with Iterative Training), an unsupervised ASR framework with a segmentation model and a phoneme prediction model. The segmentation model determines segmental structure boundaries in speech signals, while the phoneme prediction model assigns a phoneme to each segmental structure. After properly initializing the phoneme prediction model, we use an iterative algorithm that alternates between two stages to train the segmentation model and refine the phoneme prediction model. The first stage trains the segmentation model. Since we do not have the ground truth segmental boundaries, we use reinforcement learning to train the segmentation model to favor segmentations that yield phoneme sequence predictions with a lower perplexity. We experiment with various learning objectives and implement them as reward functions for learning the segmentation model. In the second stage, based on the boundaries predicted by the segmentation model learned in the previous stage, we use an adversarial loss similar to generative adversarial networks (GANs) [20] to train the phoneme prediction model.
@@
We conduct extensive experiments on LibriSpeech [45], TIMIT [19], and Multilingual LibriSpeech (MLS) [50] to compare the phoneme/phone error rate (PER) and word error rate (WER) with prior works. We show that our method outperforms all prior UASR methods on LibriSpeech, TIMIT, and five languages in MLS when using the same amount of training data. We perform thorough ablation studies to show that iterative training and the rewards we design are critical to the performance of Reborn. By analyzing the segmental structure obtained by our segmentation model, we find that the segmental structures are acoustic units smaller than phonemes, which helps the phoneme prediction model predict more accurate transcriptions. To facilitate further research and replication, our code and models are available at https://github.com/andybi7676/reborn-uasr.
2 Related Work
@@
An ASR model takes speech signals as the input and predicts the textual transcriptions. Unsupervised ASR (UASR) aims to train an ASR model without access to paired speech-text data. Instead, the only available training data is unlabeled speech data and unlabeled text data, while the correspondence between speech and text is not available. In this paper, we follow prior UASR works [33, 4] to predict phoneme transcriptions from speech signals. To achieve this, a lexicon is required to transform the unlabeled text into phoneme sequences. Learning the mapping between speech signals and phoneme sequences can be formulated as a distribution matching problem, where we want to learn an ASR model whose output phoneme sequences match the distribution of real phoneme sequences.
@@
Yeh et al. [64] address this distribution matching problem using an unsupervised loss function based on empirical output distribution matching [34], which guides the ASR model to produce phoneme sequences with statistical distributions close to real phoneme sequences. Wang et al. [60] further extend this approach by matching the $N$-skipgram and positional unigram distributions.
@@
Liu et al. [33] propose to use the generative adversarial network (GAN) [20] for UASR. The GAN solves the distribution matching problem by training a generator whose output distribution resembles a target distribution, and a discriminator is trained to distinguish the outputs of the generator and the samples from the target distribution. When using GAN to train unsupervised ASR, the generator is the phoneme prediction model. The phoneme prediction model takes in speech features of the segmental structures in the speech signal, and outputs a phoneme transcription. Prior works rely on hand-crafted or separately trained unsupervised phoneme segmentation models to find the boundary of the segmental structures [61, 33, 11].
@@
Figure 1: (a) and (b): Reborn iterates between using RL to train the segmentation model and using adversarial training to train the phoneme prediction model. (c): An illustration of the segmentation/boundary merging. 1 means the start of a segment while 0 is not. Given the original segmentation and the predicted phoneme sequence, we merge the segments that result in the same phoneme prediction into the same segment, yielding the merged boundary.
@@
In these prior methods, the segmentation is either non-trainable or learned separately from the phoneme prediction model. Reborn differs from prior works by introducing a trainable segmentation model tailored for the phoneme prediction model, and the segmentation model and phoneme prediction model can be iteratively polished. wav2vec-U 2.0 [30] simplifies the feature preprocessing step in wav2vec-U and does not explicitly consider the segmentation structure in their model. Reborn has performance better than wav2vec-U 2.0 on almost all datasets, showing that learning the segmentation boundary for feature preprocessing is important for the performance of UASR.
Although recent UASR is originally to be a cross-modality distribution matching problem, it is also related to representation learning. For example, the speech features might be extracted from self-supervised speech foundation models; and the segmentation problem is also commonly investigated in tasks like acoustic unit discovery. We discuss these two related topics further in Appendix D.
@@ 3 Method: Reborn
The difficulty of mapping speech signals to their corresponding phoneme transcriptions lies in the segmental structures in speech whose boundaries are unknown. To tackle this challenge, we propose Reborn, which trains a UASR system with unpaired speech and text data. Reborn contains a segmentation model and a phoneme prediction model. The segmentation model takes the speech feature as input and determines the boundaries of the segmental structures in the speech signal, and the phoneme prediction model predicts a phoneme for each segmental structure. We will use the terms segment and segmental structure interchangeably. In our paper, we do not use the term segment in the exact sense as in linguistics, where segment refers to discrete units that can be identified in the stream of speech [15]. We use segment to broadly refer to a span in the speech, which may be a meaningful unit (e.g., a phone or word) or a span in the speech identified by the segmentation model.
The overall training process of Reborn is outlined in Figure 1. First, we initialize the phoneme prediction model using wav2vec-U (Section 3.3). Next, the training of Reborn iterates through two stages: Stage 1 (Figure 1(a), Section 3.1): Training the segmentation model to learn a better segmentation until the segmentation model converges while fixing the phoneme prediction model. Stage 2 (Figure 1(b), Section 3.2): Training the phoneme predictor based on the segment boundaries predicted by the segmentation model until the phoneme prediction model converges. The iterative process ceases when the UASR performance does not improve over the previous iteration. In UASR, we cannot use PER as a validation metric to select the best model or determine when to stop training. We use an unsupervised evaluation metric to achieve it instead, detailed in Appendix C.4 and  C.5.
@@ 3.1.1 Segmentation Model
Given an input speech signal, we use a self-supervised feature extractor model to extract the speech features $X=[x_{1},\cdots,x_{T}]$ from the waveform. The segmentation model takes the speech features and predicts the boundary of the segmental structure in the speech features. For each feature $x_{i}$, the segmentation model assigns a binary value $\hat{b}_{i}$, indicating whether $x_{i}$ is the first frame of a segment. We split $X$ into segments based on $\hat{b}=\hat{b}_{1:T}$ and mean-pool the features within the same segment. The mean-pooled features $[s_{1},\cdots,s_{T^{\prime}}]$ are forwarded through the phoneme prediction model to obtain the phoneme prediction $[y_{1},\cdots,y_{T^{\prime}}]$ using greedy decoding. The resulting phoneme prediction will be de-duplicated into $[y^{\prime}_{1},\cdots,y^{\prime}_{M}]$ by removing the same consecutive phoneme.
@@ 3.1.2 Training the Segmentation Model with RL
To help the phoneme prediction model predict phonemes, we want to train the segmentation model to capture the segmental structure in speech. However, the optimal segmentation boundary is not available for training. Recognizing that the segmentation quality directly affects the phoneme prediction of the phoneme prediction model, we estimate the quality of the phoneme prediction to guide the segmentation model training. The phoneme prediction model is from the previous Reborn iteration or initialized from wav2vec-U in the first Reborn iteration, and it is fixed when training the segmentation model. Given that the feature segmentation based on boundary decision is inherently non-differentiable, we leverage RL to train the segmentation model. While related works may employ techniques like soft monotonic alignment or straight-through estimator to approximate gradients [7, 60], we find that RL provides a natural and suitable approach for our scenario, where the estimated quality of the phoneme sequence across different segmentations serves as the RL reward.
We train the segmentation model using the policy gradient method [57] based on the REINFORCE algorithm [63]. For each utterance, we calculate an utterance-wise reward $R$ (defined in Eq. 4 in the next subsection) from the de-duplicated phoneme prediction to train the segmentation model. Based on the policy gradient method, the segmentation model $\pi_{\theta}$ is optimized using the following gradient: $\mathbb{E}_{\hat{b}\sim\pi_{\theta}}\left[\nabla_{\theta}\log\pi_{\theta}(\hat{b}|X)R\right]$, where $\mathbb{E}$ is taken over $\hat{b}$ sampled from $\pi_{\theta}$ and approximated with the mean of a batch of training sequences.
@@ 3.1.3 Reward Designs
We introduce some notations: Given an utterance, we use the currently trained segmentation model $\pi_{\theta}$ for segmentation and the phoneme prediction model trained in the last Reborn iteration to obtain the de-duplicated phoneme prediction sequence $Y^{\prime}_{\theta}$. For the same utterance, we use the segmentation model from the previous Reborn iteration $\pi_{\theta-1}$ for features segmentation and the same phoneme prediction model to obtain another de-duplicated phoneme prediction sequence, denoted as $Y^{\prime}_{\theta-1}$. In the first Reborn iteration, $Y^{\prime}_{\theta-1}$ is the de-duplicated phoneme prediction from wav2vec-U.
Perplexity Difference Reward. The perplexity difference reward is designed to favor phoneme segmentation better than the segmentation learned in the previous iteration. Intuitively, a better phoneme segmentation prediction $\hat{b}_{1:T}$ should yield a more reasonable phoneme prediction $y^{\prime}_{1:M}$. We use perplexity (PPL), the negative likelihood of a phoneme sequence scored by a phoneme language model (LM), to evaluate how reasonable a phoneme sequence is. Perplexity measures how likely a phoneme sequence can be observed in the real world; a phoneme sequence with a lower perplexity means it is more likely to be observed in real-world phoneme datasets. We use a 4-gram phoneme LM trained on the phonemicized unlabeled text corpora. To guide the segmentation model to generate a segmentation with a better phoneme prediction than the phoneme predictions obtained in the previous iteration, we define perplexity difference reward as follows:
@@
Edit distance reward. We use Levenshtein distance $d_{\text{Lev}}$ as the edit distance to calculate the difference between $Y^{\prime}_{\theta-1}$ and $Y^{\prime}_{\theta}$. We take the normalized negative edit distance as the reward, which is defined in Eq. 2
Length difference reward. We use the length difference reward $R_{\text{len}}$ to guide the segmentation model to predict segmentation such that the length of $Y^{\prime}_{\theta}$ does not differ significantly from the length of $Y^{\prime}_{\theta-1}$. The length difference reward for an utterance is defined in Eq. 3, where $|Y^{\prime}|$ is the length of $Y^{\prime}$.
@@
where $c_{\text{ppl}}$, $c_{\text{edit}}$ and $c_{\text{len}}$ are the weighting coefficients. During training, $R_{\text{ppl}}$, $R_{\text{edit}}$ and $R_{\text{len}}$ are normalized within each batch. Appendix C.5 discusses how the coefficients are determined.
@@ 3.1.4 Initializing $\pi_{\theta}$ with Behavior Cloning
Before training the segmentation model $\pi_{\theta}$ with RL, we initialize it with behavior cloning (BC) [47, 48]. BC uses the supervised objective to train the segmentation model to predict the boundaries in speech features. Given speech features $x_{1:T}$, the segmentation model is trained to predict the 0/1 boundary $\hat{b}_{i}$ by using some boundaries $b_{i}$ as the target. In Reborn, the target boundary for BC is the merged boundary (Section 3.1.5) obtained using $\pi_{\theta-1}$, the segmentation model learned in the previous iteration. We then frame BC as a binary classification task and optimize it with the cross-entropy loss. Note that in the first Reborn iteration, we use the $k$-means-based boundary from wav2vec-U as the BC target.
@@ 3.2 Stage 2: Training the Phoneme Prediction Model
After obtaining the merged phoneme boundaries, we train the phoneme predictor model using GAN training. The phoneme prediction model is the generator in GAN training. The generator aims to output phoneme predictions that look like real phoneme sequences to fool the discriminator. The discriminator takes in a phoneme sequence, which can be the output of the generator or a phoneme sequence from the unpaired text corpora, and the goal of the discriminator is to distinguish whether the input phoneme sequences are outputs of the generator. The generator and discriminator are updated using the loss in GAN training. In this stage, the parameters of the segmentation model are not updated, and the generator is initialized from the previous iteration. We discuss the effect of applying boundary merging and parameter initialization in Appendix A.1